Determine the final temperature of the liquid when you mix:

m1 = 1.00 kg water at t1 = 10.0 °C

and

m2 = 2.00 kg water at t2 = 40.0 °C.

Assume that no heat is lost to the environment.

A numerical value is required for credit.

SHOW YOUR WORK.


Formulas: Heat: Q = mcAT Q1 + Q2 + ...=0 Calorimetry:
Answer #1

Hot balancing heat transfer a cold +mzcp (tz-T) = mi cp (T-t) - 2 (yo-T) = 1(T-10) T= 30°c